OpenStudy (anonymous):

Solve 5c + 4 = -26. 6 -6 -4.4 3

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

All I know is that I have to make C by its self but I forgot how to do that

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (fibonaccichick666):

how do you undo addition?

OpenStudy (anonymous):

subtraction?

OpenStudy (fibonaccichick666):

yup, so what will you subtract from both sides to isolate 5c?

OpenStudy (helpblahblahblah):

if im not mistaking you do -26-4 which equals -30 then you come with 5c=-30 so then you divide and that is -6 but however i dont know if i did it right lol sorry

OpenStudy (anonymous):

5-5 would cancel each other out-then you would have to do -26-5=-31

OpenStudy (fibonaccichick666):

no, not yet

OpenStudy (fibonaccichick666):

you need to get rid of the 4 to start

OpenStudy (anonymous):

:P how??

OpenStudy (fibonaccichick666):

again, how do you undo addition?

OpenStudy (anonymous):

subtraction

OpenStudy (anonymous):

help me @bibby

OpenStudy (fibonaccichick666):

so, if you want to rid yourself of the plus 4 how can you undo it?

OpenStudy (anonymous):

@FibonacciChick666 do I subtract 5-4??

OpenStudy (bibby):

you're being helped, this literally hurts me to read

OpenStudy (bibby):

that's not proper english, you break my brain

OpenStudy (fibonaccichick666):

no, you have a 5c, not a 5, since you do not have like terms you cannot combine the 5 and 4 in any way shape or form. please read this page before i continue so I don't lose my mind http://www.purplemath.com/modules/solvelin.htm

OpenStudy (anonymous):

do I do -26-4?

OpenStudy (fibonaccichick666):

please write out the full step not just one side

OpenStudy (anonymous):

ahhh I don't know what you mean!!!

OpenStudy (fibonaccichick666):

did you read the page like I asked?

OpenStudy (fibonaccichick666):

How did they show their work? do it the same way

OpenStudy (anonymous):

\[4-5c=1c \] \[4-26=either -30 or -22\] right??

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

@bibby am I doing this right?

OpenStudy (fibonaccichick666):

what? where did you get 4-5c from?

OpenStudy (fibonaccichick666):

you have \[5c+4=-26\]

OpenStudy (anonymous):

thats what purple math said to do

OpenStudy (fibonaccichick666):

no, it didn't please look about half way down to where the example with x+6= something is

OpenStudy (anonymous):

the answer is -6 right

OpenStudy (fibonaccichick666):

I'm not telling you what the answer is or verifying until you show me proper work

OpenStudy (anonymous):

wtf man I'm just asking if its -6

OpenStudy (anonymous):

its not my fault I don't know wtf I'm doing

OpenStudy (fibonaccichick666):

it is your fault, you have not practiced enough yet

OpenStudy (fibonaccichick666):

and I reiterate: I'm not telling you what the answer is or verifying until you show me proper work

OpenStudy (anonymous):

I'm trying to do the work and its just more confusing if your not helping me

OpenStudy (bibby):

why don't you just write out the work

OpenStudy (anonymous):

i did and i'm still confused this isn't helping!!

OpenStudy (bibby):

write it out. for example 2c + 8 = 4 -8 = -8 --------------- 2c = -4

OpenStudy (anonymous):

i need like some one explaining how to do it not yelling at me for proper work

OpenStudy (anonymous):

thats what I did and this chick said it was wrong

OpenStudy (bibby):

I have no idea what you did, but it's certainly not what she asked you to do

OpenStudy (anonymous):

well mine trying ok gosh

OpenStudy (anonymous):

wait don't to negatives equal a positive?

OpenStudy (anonymous):

damn it I can't even get the right answer I did what you did and now I can't get the freakn answer!

OpenStudy (anonymous):

Why do I suck at math!!!

OpenStudy (bibby):

You have the right answer

OpenStudy (anonymous):

what -6 because thats not even my answer thats the girl @HelpBlahBlahBlah answer and she's smarter then me because I'm in algebra 1 and she's going into pre algebra and she probably got it right!!

OpenStudy (bibby):

why don't you understand that getting the answer from random shmucks on the internet won't teach you as well as doing problems will? 5c + 4 = -26. - 4 = - 4 ------------ 5c = -30

OpenStudy (anonymous):

Dude I understand that ok its just I can't do the freaking math because I'm a freaking idiot that sucks at everything besides any kind of science or sports! I mean I can't even subtract 2 negative numbers!

OpenStudy (anonymous):

and -30 isn't even the answer!

OpenStudy (bibby):

5c = -30 they're asking for c...

OpenStudy (anonymous):

-6...

OpenStudy (anonymous):

i don't know I did the math and probably wrong just because I divided -30/5=-6

OpenStudy (bibby):

that's what you were supposed to do, yeah well at least step 2

OpenStudy (anonymous):

I know what to I just don't know how to do it 100% correctly!